# Mock Observations of Multiple Stellar Populations in Tidal Streams of Palomar 5 for the Chinese Space Station Survey Telescope

### 13 Mar 2026 v1

@@ II.1 N-body Simulation Data
Our simulation data are derived from the previous work on Palomar 5 by Wang et al. (2023). The authors conducted simulations of Palomar 5-like clusters using the high-performance $N$-body code petar (Wang et al., 2020a). This code can efficiently model the evolution of star clusters and tidal streams. It uses the high-performance parallel framework fdps (Iwasawa et al., 2016, 2020), which implements the hybrid $P^{3}T$ method (Oshino et al., 2011): a high-precision direct $N$-body integrator (4th-order Hermite with slowdown-algorithmic regularization (SDAR) method (Wang et al., 2020c)) for short-range interactions and multiple systems, coupled with a particle-tree method (Barnes and Hut, 1986) for long-range forces. The code also includes updated stellar evolution codes (sse/bse (Hurley et al., 2000, 2002; Banerjee et al., 2020)) for evolving individual stars and binaries, and supports external galaxy potential through galpy (Bovy, 2015). These features make petar essential for modelling star-by-star dynamical evolution of GCs with tidal tails.
@@ II.3 Synthetic Spectra
The absolute flux spectrum serves two key purposes: (1) to serve as input for the CSST/SC mock observation pipeline, and (2) to derive both absolute and apparent magnitudes. We used the package spectrum (version 2.77c) to synthesize the spectra. spectrum performs a standard computation assuming local thermodynamic equilibrium (LTE) and a plane-parallel atmosphere. We used atlas9 (Castelli and Kurucz, 2004) as input stellar atmosphere models. These models are based on the solar abundances by Grevesse and Sauval (1998). The model grids are computed for $T_{\mathrm{eff}}$ from 3,500 K to 50,000 K, $\log{g}$ from 0.0 dex to 5.0 dex, and [M/H] = +0.5, +0.2, 0.0, -0.5, -1.0, -1.5, -2.0, -2.5, -3.5, -4, and -5.5. The input line lists were taken from the spectrum package website, and cover a wavelength range of 900 Å–40000 Å. According to the wavelength range of the CSST, we computed the synthetic spectra from 2500 to 11000 Å.
@@ III.1 Photometric Result
In the study of helium variations in star clusters using CSST/SC, Ji et al. (2023) demonstrated that 2P stars, driven by He-enrichment, are bluer than 1P stars across all color bands. The study further highlighted that the unique CSST/SC $NUV$, $u$ passbands are particularly effective in maximizing the color separation between helium-normal and helium-enriched populations. According to Milone et al. (2015b, a), N-enriched stars are also CO-depleted, with key CNO-related spectral features predominantly distributed in the UV to blue wavelength range. Therefore, we expect that the $NUV$ and $u$ bands of CSST/SC will resolve a larger color difference between the 1P and 2P in both MS and RGB phase. Figure 5 presents a series of CMDs in the $g$-band versus various colors to characterize the stellar populations in region 1. We selected the $i$-band as the subtrahend to ensure a long wavelength baseline, thereby emphasizing colors involving UV bands. Data points are color-coded based on known labels (top two rows of Figure 5): red for 1P, blue for 2P, and gray for field stars. The figure reveals that while CMDs utilizing CSST/SC UV photometry are powerful in disentangling stellar populations on the giant branch, the $g$ vs. gig-i CMD provides a clearer distinction between the 1P and 2P on the MS phase at a 150 s exposure. This result appears inconsistent with previous findings by Ji et al. (2023), who reported diminishing color differences in both the MS and RGB as the minuend filter changes from NUV to i. To further investigate this discrepancy, we turn to bottom two rows of Figure 5, which presents normalized stellar number density maps constructed without prior knowledge of stellar labels from region 1. These maps corroborate that the $g$-band provides superior separation for the MS phase, whereas on the giant branch, the member stars blend with the field stars, showing no clear separation. Notably, the $NUV$ and $u$ bands fail to reveal clear distinctions between the different populations.
@@
This phenomenon can be attributed to several factors. First, He-enrichment causes 2P stars to become bluer even in the $g$ band. Second, as reviewed by Zhan (2021), the system throughput efficiency in the $g$ band is significantly higher than that in the $NUV$ and $u$ bands, allowing better photometric precision. Figure 6 shows that at a comparable magnitude level, the photometric errors in the $g$-band (green points) are systematically and significantly lower than those in the $NUV$ (blue points) and the $u$-band (orange points). This improved precision produces tighter stellar sequences in the CMD, thereby amplifying the subtle color differences between distinct stellar populations and enabling clearer resolution.
Additionally, the CMDs for region 1 exhibit relatively few field stars in the MS phase compared to the giant phase. This scarcity further challenges the detection of significant color differences in the giant star phase using $NUV$ and $u$ bands during 150-second exposures without prior stellar labels. Regarding the paucity of field stars in the MS phase, we have re-examined the complete Palomar 5 stellar stream dataset and confirmed that this characteristic persists in the region 1 and region 2 (see the APPENDIX). The most probable reason is the dominance of member stars over field stars at an equal distance modulus.
Figure 6: Magnitude error as function of magnitude in region 1 for the $NUV$ (blue points), $u$ (orange points), and $g$ (green) bands.
We performed similar analyses for regions 2 and 3 in Figures 7 and 8. The photometric precision results of region 2 and region 3 are similar to Figure 6. The figure is not repeated here. Since regions 2 and 3 are located within the stellar stream of Palomar 5, the number of member stars has significantly decreased. Consequently, even with known stellar types (1P, 2P or field stars), the $NUV$ and $u$ bands can no longer clearly separate the 1P and 2P in the CMDs of Figures 7 and 8, whereas the $g$ band retains this capability. Furthermore, as observed in Figures 5, 7, and 8, the ability to disentangle stellar populations weakens in the $NUV$, $u$, and $g$ bands as the number of member stars decreases. The same conclusion can be drawn from the CMDs in Figures 7 and 8, for which no stellar types are provided.
@@ III.2 Extent of Stellar Population Contamination
The distance and faintness of the Palomar 5 stream make it particularly challenging to cleanly isolate its member stars from the substantial foreground and background contamination along the line of sight. This is further exacerbated by various factors contributing to the photometric error, such as the photometric noise, unresolved binaries, blending, PSF fitting residuals, etc. These considerations motivate us to evaluate membership assignment in which we combine the CMD information described above with $NUV$, $u$ and $g$ bands photometry. We have presented a rough method for tagging member stars (1P or 2P) and field stars in all three regions. Based on the demonstrated capability of the $NUV$, $u$, and $g$ photometric bands to distinguish MPs (as detailed in Section III.1), we constructed theoretical CMDs in the $g$ versus NUViNUV-i, uiu-i, and gig-i colors using the theoretical apparent magnitudes described in Section II.4. We then calculated the identification accuracy of member stars under the judgment of these three theoretical CMDs, respectively. The photometric magnitudes of a star that does not conform to either the 1P or 2P theoretical CMD sequences within photometric error is classified as a field star; otherwise, it is considered a member star. In specific cases where a star is consistent with both sequences, its classification (1P or 2P) is determined by comparing its distance to each fiducial sequence. Table 2 presents the classification results of member stars and field stars for the three regions. In the parentheses in Table 2, we include the contamination rate and completeness rate for each category, defined as:
@@
From Table 2, classifications using the NUViNUV-i color achieve near-perfect purity (contamination rate $\approx 0\%$ for member stars in all regions), but at the cost of very low completeness ($\sim 2\%$). This suggests that while the few stars selected by NUViNUV-i are highly reliable members, the $NUV$ band’s faintness and associated large photometric errors cause the vast majority of genuine members to be missed. In contrast, the gig-i color offers a more balanced performance. It maintains excellent purity (contamination rate $1\%$ in region 1) while achieving significantly higher completeness ($\sim 31\%$ in region 1), making it the most effective single color for member identification. The uiu-i color performs intermediately. Furthermore, the efficacy of all color indices deteriorates in stream regions (regions 2 and 3), where the field-to-member ratio is extreme. For instance, the contamination rate for (gig-i)-selected members in region 3 rises to $\sim 11\%$. This indicates that in the sparse stellar stream, CMDs alone are insufficient for reliable member discrimination due to the overwhelming field star contamination. Consequently, robust member identification in these regions will necessitate the integration of additional parameters, such as deep photometry, proper motions, and radial velocities.
@@ III.3 Deep Photometry for Palomar 5
CSST will conduct a 10-year survey, performing multi-color photometric imaging and slitless spectroscopy over 17,500 $\deg^{2}$ in wide-field and 400 $\deg^{2}$ in deep-field modes. The CSST/SC is responsible for these primary observational tasks. For the wide-field survey, each sky area will be covered at least twice in each photometric band, with a single exposure of $\approx$150 s. To detect MPs of MS and RGB phases in the $g$ band, and to ensure a SNR of no less than 100 in the $NUV$ and $u$ bands, we used the online CSST exposure time calculator (ETC) to estimate that the required exposure time needs to be at least 150s $\times$ 10 (1500 s). Therefore, we stacked ten 150 s exposures from three regions using SWarp (Bertin et al., 2002). This stacking procedure will allow us to assess the CSST’s ability to resolve different stellar populations under deep photometric conditions. Based on the previous analysis, we only discuss the deep photometry of the $NUV$, $u$, and $g$ bands.
@@
Figures 9, 10, and 11 present the results of deep photometry for regions 1, 2, and 3, respectively. With the exception of the sparse regions of the stream, the left panels demonstrate that, if member stars can be robustly identified, deep photometry in all three bands can, in principle, distinguish MPs in both the MS and RGB phases. The $g$-band reveals a more pronounced separation between the 1P and 2P MS phase compared to a single exposure of 150 s. In the cases of unknown classifications, we found that even after stacking ten exposures in the $NUV$ band (which improves the SNR by approximately a factor of 3.16), no clear distinction is observed between the different stellar populations on the MS. Furthermore, the stacked data in all bands fail to reveal MPs in the RGB phase.
@@
The difficulty in resolving MPs in extended streams like Palomar 5 directly from CMDs underscores the necessity of proper motion data. While photometric surveys from the HST and the upcoming CSST/SC provide valuable color-magnitude information, precise kinematic measurements remain essential to disentangle stream members from foreground and background contamination. Starkman et al. (2020) used proper motion data from Gaia data release 2 (Gaia DR2) to select members of the Palomar 5 stream. Their analysis showed that the proper motion distribution of stars within two projected tidal radii of Palomar 5 is sharply peaked around -2.7 $\pm$ 0.1 (mas yr-1) (Starkman et al., 2020). Based on this peak, they adopted a circular region with a radius of 3 mas yr-1 centered at ($\mu_{\alpha^{*}}$, $\mu_{\delta}$) == (-2.5, -2.5) mas yr-1 as a proper motion selection window. This window efficiently removed 81.9% of field stars. However, due to the distance of Palomar 5, its main sequence turn-off stars lie near Gaia’s detection limit ($G$ $\approx$ 21 mag). Even with the improved precision of Gaia data release 3 (Gaia DR3), the bulk of the stream’s stellar population remains not well resolved in Gaia data (Bonaca and Price-Whelan, 2025). In Gaia DR2, the proper motion uncertainty is approximately 0.07, 0.2, 1.2, and 3 mas yr-1 at $G$ << 15 mag, $G$ == 17, 20, and 21 mag, respectively (Gaia Collaboration et al., 2018). CSST will also possess a strong astrometric capability. Under ideal conditions, CSST is expected to achieve proper motion accuracies << 1 mas yr-1 for stars between 18-22 mag in $g$ band, and >> 1 mas yr-1 for those at 22-26 mag in $g$ band (Fu et al., 2023). By combining proper motion data from Gaia and CSST, we can identify member stars in both the MS and RGB phases. Once CSST is fully operational, we plan to observe the Palomar 5 cluster center and both high- and low-density regions of its tidal tails. By conducting observations across ten different epochs in multiple bands ($NUV$, $u$, $g$, ii), we aim to obtain the deep photometry and proper motion data necessary for a definitive MP analysis.
@@
Finally, based on Figure 6, a rough estimate suggests that to achieve comparable photometric precision in $NUV$ band as in the $g$-band for detecting MPs under 150 s exposures, the magnitude difference is approximately 2. Using the formula $m_{1}-m_{2}=5\log_{10}{\left(\frac{d_{1}}{d_{2}}\right)}$, we estimate that the observed stream would need to be located at a distance of $\lesssim$ 8 kpc.
@@ V Conclusion
Our simulations under 150 s exposures identify the $g$-band as the optimal filter for distinguishing 1P and 2P populations on the MS, a direct result of its superior photometric precision. Although the $NUV$ and $u$-bands cover critical spectral features, their lower photometric accuracy renders them impractical for resolving the MPs. For RGB stars, population separation remains challenging across all bands due to blending with field stars, the issue exacerbated in sparse stream regions by declining member star counts. We further evaluated member star identification using theoretical CMDs, revealing that the CSST/SC $g$-band provides the optimal balance between contamination rate and completeness rate in Palomar 5’s core. However, this band fails in the stream due to severe field star contamination, unequivocally highlighting the necessity of other methods for robust membership determination.
To improve the SNR, we simulated ten exposures of 150 s for all three regions and subsequently stacked the simulation images. Even after stacking $NUV$ band images to improve the SNR, distinct separation between different stellar populations on the MS remains unachievable. Furthermore, the stacked data in all bands failed to detect any MP signals among the RGB. Our results highlight that even with CSST’s high UV sensitivity and large FoV, the detection of MPs in distant, extended streams is limited in the absence of proper motion data. Future CSST observations of Palomar 5 and its tidal tails will employ ten different epochs in multiple bands ($NUV$, $u$, $g$, ii) to provide the deep photometry (26 mag in $g$ band) and proper motions (uncertainty: << 1 mas yr-1 at $g$ = 18-22 mag and >> 1 mas yr-1 at $g$ = 22-26 mag) necessary for a definitive MP analysis.

### 18 Mar 2026 v2

@@ II.1 N-body Simulation Data
Our simulation data are derived from the previous work on Palomar 5 by Wang et al. (2023). The authors conducted simulations of Palomar 5-like clusters using the high-performance NN-body code petar (Wang et al., 2020a). This code can efficiently model the evolution of star clusters and tidal streams. It uses the high-performance parallel framework fdps (Iwasawa et al., 2016, 2020), which implements the hybrid $P^{3}T$ method (Oshino et al., 2011): a high-precision direct NN-body integrator (4th-order Hermite with slowdown-algorithmic regularization (SDAR) method (Wang et al., 2020c)) for short-range interactions and multiple systems, coupled with a particle-tree method (Barnes and Hut, 1986) for long-range forces. The code also includes updated stellar evolution codes (sse/bse (Hurley et al., 2000, 2002; Banerjee et al., 2020)) for evolving individual stars and binaries, and supports external galaxy potential through galpy (Bovy, 2015). These features make petar essential for modelling star-by-star dynamical evolution of GCs with tidal tails.
@@ II.3 Synthetic Spectra
The absolute flux spectrum serves two key purposes: (1) to serve as input for the CSST/SC mock observation pipeline, and (2) to derive both absolute and apparent magnitudes. We used the package spectrum (version 2.77c) to synthesize the spectra. spectrum performs a standard computation assuming local thermodynamic equilibrium (LTE) and a plane-parallel atmosphere. We used atlas9 (Castelli and Kurucz, 2004) as input stellar atmosphere models. These models are based on the solar abundances by Grevesse and Sauval (1998). The model grids are computed for $T_{\mathrm{eff}}$ from 3,500 K to 50,000 K, $\log{g}$ from 0.0 dex to 5.0 dex, and [M/H] = +0.5, +0.2, 0.0, -0.5, -1.0, -1.5, -2.0, -2.5, -3.5, -4, and -$5.$ The input line lists were taken from the spectrum package website, and cover a wavelength range of 900 Å–40000 Å. According to the wavelength range of the CSST, we computed the synthetic spectra from 2500 to 11000 Å.
@@ III.1 Photometric Result
In the study of helium variations in star clusters using CSST/SC, Ji et al. (2023) demonstrated that 2P stars, driven by He-enrichment, are bluer than 1P stars across all color bands. The study further highlighted that the unique CSST/SC NUVNUV, $u$ passbands are particularly effective in maximizing the color separation between helium-normal and helium-enriched populations. According to Milone et al. (2015b, a), N-enriched stars are also CO-depleted, with key CNO-related spectral features predominantly distributed in the UV to blue wavelength range. Therefore, we expect that the NUVNUV and $u$ bands of CSST/SC will resolve a larger color difference between the 1P and 2P in both MS and RGB phase. Figure 5 presents a series of CMDs in the $g$-band versus various colors to characterize the stellar populations in region 1. We selected the $i$-band as the subtrahend to ensure a long wavelength baseline, thereby emphasizing colors involving UV bands. Data points are color-coded based on known labels (top two rows of Figure 5): red for 1P, blue for 2P, and gray for field stars. The figure reveals that while CMDs utilizing CSST/SC UV photometry are powerful in disentangling stellar populations on the giant branch, the $g$ vs. gig-i CMD provides a clearer distinction between the 1P and 2P on the MS phase at a 150 s exposure. This result appears inconsistent with previous findings by Ji et al. (2023), who reported diminishing color differences in both the MS and RGB as the minuend filter changes from NUV to i. To further investigate this discrepancy, we turn to bottom two rows of Figure 5, which presents normalized stellar number density maps constructed without prior knowledge of stellar labels from region 1. These maps corroborate that the $g$-band provides superior separation for the MS phase, whereas on the giant branch, the member stars blend with the field stars, showing no clear separation. Notably, the NUVNUV and $u$ bands fail to reveal clear distinctions between the different populations.
@@
This phenomenon can be attributed to several factors. First, He-enrichment causes 2P stars to become bluer even in the $g$ band. Second, as reviewed by Zhan (2021), the system throughput efficiency in the $g$ band is significantly higher than that in the NUVNUV and $u$ bands, allowing better photometric precision. Figure 6 shows that at a comparable magnitude level, the photometric errors in the $g$-band (green points) are systematically and significantly lower than those in the NUVNUV (blue points) and the $u$-band (orange points). This improved precision produces tighter stellar sequences in the CMD, thereby amplifying the subtle color differences between distinct stellar populations and enabling clearer resolution.
Additionally, the CMDs for region 1 exhibit relatively few field stars in the MS phase compared to the giant phase. This scarcity further challenges the detection of significant color differences in the giant star phase using NUVNUV and $u$ bands during 150-second exposures without prior stellar labels. Regarding the paucity of field stars in the MS phase, we have re-examined the complete Palomar 5 stellar stream dataset and confirmed that this characteristic persists in the region 1 and region 2 (see the APPENDIX). The most probable reason is the dominance of member stars over field stars at an equal distance modulus.
Figure 6: Magnitude error as function of magnitude in region 1 for the NUVNUV (blue points), $u$ (orange points), and $g$ (green) bands.
We performed similar analyses for regions 2 and 3 in Figures 7 and 8. The photometric precision results of region 2 and region 3 are similar to Figure 6. The figure is not repeated here. Since regions 2 and 3 are located within the stellar stream of Palomar 5, the number of member stars has significantly decreased. Consequently, even with known stellar types (1P, 2P or field stars), the NUVNUV and $u$ bands can no longer clearly separate the 1P and 2P in the CMDs of Figures 7 and 8, whereas the $g$ band retains this capability. Furthermore, as observed in Figures 5, 7, and 8, the ability to disentangle stellar populations weakens in the NUVNUV, $u$, and $g$ bands as the number of member stars decreases. The same conclusion can be drawn from the CMDs in Figures 7 and 8, for which no stellar types are provided.
@@ III.2 Extent of Stellar Population Contamination
The distance and faintness of the Palomar 5 stream make it particularly challenging to cleanly isolate its member stars from the substantial foreground and background contamination along the line of sight. This is further exacerbated by various factors contributing to the photometric error, such as the photometric noise, unresolved binaries, blending, PSF fitting residuals, etc. These considerations motivate us to evaluate membership assignment in which we combine the CMD information described above with NUVNUV, $u$ and $g$ bands photometry. We have presented a rough method for tagging member stars (1P or 2P) and field stars in all three regions. Based on the demonstrated capability of the NUVNUV, $u$, and $g$ photometric bands to distinguish MPs (as detailed in Section III.1), we constructed theoretical CMDs in the $g$ versus NUViNUV-i, uiu-i, and gig-i colors using the theoretical apparent magnitudes described in Section II.4. We then calculated the identification accuracy of member stars under the judgment of these three theoretical CMDs, respectively. The photometric magnitudes of a star that does not conform to either the 1P or 2P theoretical CMD sequences within photometric error is classified as a field star; otherwise, it is considered a member star. In specific cases where a star is consistent with both sequences, its classification (1P or 2P) is determined by comparing its distance to each fiducial sequence. Table 2 presents the classification results of member stars and field stars for the three regions. In the parentheses in Table 2, we include the contamination rate and completeness rate for each category, defined as:
@@
From Table 2, classifications using the NUViNUV-i color achieve near-perfect purity (contamination rate $\approx 0\%$ for member stars in all regions), but at the cost of very low completeness ($\sim 2\%$). This suggests that while the few stars selected by NUViNUV-i are highly reliable members, the NUVNUV band’s faintness and associated large photometric errors cause the vast majority of genuine members to be missed. In contrast, the gig-i color offers a more balanced performance. It maintains excellent purity (contamination rate $1\%$ in region 1) while achieving significantly higher completeness ($\sim 31\%$ in region 1), making it the most effective single color for member identification. The uiu-i color performs intermediately. Furthermore, the efficacy of all color indices deteriorates in stream regions (regions 2 and 3), where the field-to-member ratio is extreme. For instance, the contamination rate for (gig-i)-selected members in region 3 rises to $\sim 11\%$. This indicates that in the sparse stellar stream, CMDs alone are insufficient for reliable member discrimination due to the overwhelming field star contamination. Consequently, robust member identification in these regions will necessitate the integration of additional parameters, such as deep photometry, proper motions, and radial velocities.
@@ III.3 Deep Photometry for Palomar 5
CSST will conduct a 10-year survey, performing multi-color photometric imaging and slitless spectroscopy over 17,500 $\deg^{2}$ in wide-field and 400 $\deg^{2}$ in deep-field modes. The CSST/SC is responsible for these primary observational tasks. For the wide-field survey, each sky area will be covered at least twice in each photometric band, with a single exposure of $\approx$150 s. To detect MPs of MS and RGB phases in the $g$ band, and to ensure a SNR of no less than 100 in the NUVNUV and $u$ bands, we used the online CSST exposure time calculator (ETC) to estimate that the required exposure time needs to be at least 150s $\times$ 10 (1500 s). Therefore, we stacked ten 150 s exposures from three regions using SWarp (Bertin et al., 2002). This stacking procedure will allow us to assess the CSST’s ability to resolve different stellar populations under deep photometric conditions. Based on the previous analysis, we only discuss the deep photometry of the NUVNUV, $u$, and $g$ bands.
@@
Figures 9, 10, and 11 present the results of deep photometry for regions 1, 2, and 3, respectively. With the exception of the sparse regions of the stream, the left panels demonstrate that, if member stars can be robustly identified, deep photometry in all three bands can, in principle, distinguish MPs in both the MS and RGB phases. The $g$-band reveals a more pronounced separation between the 1P and 2P MS phase compared to a single exposure of 150 s. In the cases of unknown classifications, we found that even after stacking ten exposures in the NUVNUV band (which improves the SNR by approximately a factor of 3.16), no clear distinction is observed between the different stellar populations on the MS. Furthermore, the stacked data in all bands fail to reveal MPs in the RGB phase.
@@
The difficulty in resolving MPs in extended streams like Palomar 5 directly from CMDs underscores the necessity of proper motion data. While photometric surveys from the HST and the upcoming CSST/SC provide valuable color-magnitude information, precise kinematic measurements remain essential to disentangle stream members from foreground and background contamination. Starkman et al. (2020) used proper motion data from Gaia data release 2 (Gaia DR2) to select members of the Palomar 5 stream. Their analysis showed that the proper motion distribution of stars within two projected tidal radii of Palomar 5 is sharply peaked around -2.7 $\pm$ 0.1 (mas yr-1) (Starkman et al., 2020). Based on this peak, they adopted a circular region with a radius of 3 mas yr-1 centered at ($\mu_{\alpha^{*}}$, $\mu_{\delta}$) == (-2.5, -2.5) mas yr-1 as a proper motion selection window. This window efficiently removed 81.9% of field stars. However, due to the distance of Palomar 5, its main sequence turn-off stars lie near Gaia’s detection limit ($G$ $\approx$ 21 mag). Even with the improved precision of Gaia data release 3 (Gaia DR3), the bulk of the stream’s stellar population remains not well resolved in Gaia data (Bonaca and Price-Whelan, 2025). In Gaia DR2, the proper motion uncertainty is approximately 0.07, 0.2, 1.2, and 3 mas yr-1 at $G$ << 15 mag, $G$ == 17, 20, and 21 mag, respectively (Gaia Collaboration et al., 2018). CSST will also possess a strong astrometric capability. Under ideal conditions, CSST is expected to achieve proper motion accuracies << 1 mas yr-1 for stars between 18-22 mag in $g$ band, and >> 1 mas yr-1 for those at 22-26 mag in $g$ band (Fu et al., 2023). By combining proper motion data from Gaia and CSST, we can identify member stars in both the MS and RGB phases. Once CSST is fully operational, we plan to observe the Palomar 5 cluster center and both high- and low-density regions of its tidal tails. By conducting observations across ten different epochs in multiple bands (NUVNUV, $u$, $g$, ii), we aim to obtain the deep photometry and proper motion data necessary for a definitive MP analysis.
@@
Finally, based on Figure 6, a rough estimate suggests that to achieve comparable photometric precision in NUVNUV band as in the $g$-band for detecting MPs under 150 s exposures, the magnitude difference is approximately 2. Using the formula $m_{1}-m_{2}=5\log_{10}{\left(\frac{d_{1}}{d_{2}}\right)}$, we estimate that the observed stream would need to be located at a distance of $\lesssim$ 8 kpc.
@@ V Conclusion
Our simulations under 150 s exposures identify the $g$-band as the optimal filter for distinguishing 1P and 2P populations on the MS, a direct result of its superior photometric precision. Although the NUVNUV and $u$-bands cover critical spectral features, their lower photometric accuracy renders them impractical for resolving the MPs. For RGB stars, population separation remains challenging across all bands due to blending with field stars, the issue exacerbated in sparse stream regions by declining member star counts. We further evaluated member star identification using theoretical CMDs, revealing that the CSST/SC $g$-band provides the optimal balance between contamination rate and completeness rate in Palomar 5’s core. However, this band fails in the stream due to severe field star contamination, unequivocally highlighting the necessity of other methods for robust membership determination.
To improve the SNR, we simulated ten exposures of 150 s for all three regions and subsequently stacked the simulation images. Even after stacking NUVNUV band images to improve the SNR, distinct separation between different stellar populations on the MS remains unachievable. Furthermore, the stacked data in all bands failed to detect any MP signals among the RGB. Our results highlight that even with CSST’s high UV sensitivity and large FoV, the detection of MPs in distant, extended streams is limited in the absence of proper motion data. Future CSST observations of Palomar 5 and its tidal tails will employ ten different epochs in multiple bands (NUVNUV, $u$, $g$, ii) to provide the deep photometry (26 mag in $g$ band) and proper motions (uncertainty: << 1 mas yr-1 at $g$ = 18-22 mag and >> 1 mas yr-1 at $g$ = 22-26 mag) necessary for a definitive MP analysis.